# Exact calculation of degrees for lattice equations: a singularity approach

### v1

@@ 6. Conclusions and Discussion
In this paper, we proposed a new method to rigorously calculate exact degrees for lattice equations. Our basic idea was to extend Halburd’s method to lattice equations and to gave a framework to make the calculations rigorous. If an equation satisfies some conditions, we can compute the degrees from its singularity patterns by solving a system of linear recurrences.
@@
It should be mentioned that computing the degrees for a quad equation by deriving a linear partial difference equation satisfied by the degrees has already proposed in previous studies [40, 41]. There, one needs to assume some conjecture to ensure the calculations are rigorous. Unfortunately, our theory does not give a direct answer to their conjecture since the degrees considered in their papers are total degrees while the degrees in this paper are individual degrees. However, there is expected to be a strong connection between the strategy used in the previous studies and our theory.
@@
First, let us consider whether our theory can be applied to other types of lattice equations, such as non-quad equations on a two dimensional lattice or equations on a higher-dimensional lattice. Apparently, our main idea is still effective and many lemmas and propositions, such as Lemma 4.4 and Propositions 3.19, 3.22, are still valid in general cases, with appropriate modifications. However, the proofs of some lemmas, such as Lemmas 4.9 and 4.10, are not valid for completely general lattice equations. It seems that how difficult it is to show these lemmas varies depending on the configuration of the shift vectors appearing in a lattice equation. In the case of quad equations, the polytope obtained as the convex hull of the shift vectors appearing in an equation, i.e., $\operatorname{conv}\{(0,0),(-1,0),(0,-1),(-1,-1)\}$, is a unit square. The polytope has two edges attached to the origin $(0,0)$, but neither of them has a shift vector in the middle. This property played an essential role in the proofs of Lemmas 4.9 and 4.10. Therefore, to prove these lemmas for general lattice equations would require some knowledge of convex geometry.
@@ Appendix A Essential choices of domain
Appendix B Substitution and derivative
In this appendix, let $K$ be an arbitrary field of characteristic 00 and consider the rational function field $K(z,w)$. We allow $\infty$ as the result of substitution throughout this paper. This is why we need some unusual definitions and lemmas about substitution and derivative.
@@ Appendix B Substitution and derivative
Next, we show the converse. Let $\partial_{w}f(z^{*},w)=0$. If $f(z^{*},w)=\infty$, then it is obviously independent of $w$. Therefore, we may assume that $f(z^{*},w)\neq\infty$. In this case, usual properties of derivative hold and thus $f(z^{*},w)$ does not depend on $w$. If $\partial_{w}\frac{1}{f(z^{*},w)}=0$, one can show in the same way that $g(z^{*},w)$ does not depend on $w$ where $g(z,w)=\frac{1}{f(z,w)}$. ∎

### v2

@@ 6. Conclusions and Discussion
In this paper, we proposed a new method to rigorously calculate exact degrees for lattice equations. Our basic idea was to extend Halburd’s method to lattice equations and to give a framework to make the calculations rigorous. If an equation satisfies some conditions, we can compute the degrees from its singularity patterns by solving a system of linear recurrences.
@@
It should be mentioned that computing the degrees for a quad equation by deriving a linear partial difference equation satisfied by the degrees has already been proposed in previous studies [40, 41]. There, one needs to assume some conjecture to ensure the calculations are rigorous. Unfortunately, our theory does not give a direct answer to their conjecture since the degrees considered in their papers are total degrees while the degrees in this paper are individual degrees. However, there is expected to be a strong connection between the strategy used in the previous studies and our theory.
@@
First, let us consider whether our theory can be applied to other types of lattice equations, such as non-quad equations on a two-dimensional lattice or equations on a higher-dimensional lattice. Apparently, our main idea is still effective. In fact, many lemmas and propositions, such as Lemma 4.4 and Propositions 3.19, 3.22, are still valid in general cases, with appropriate modifications. However, the proofs of some lemmas, such as Lemmas 4.9 and 4.10, are not valid for completely general lattice equations. It seems that how difficult it is to show these lemmas varies depending on the configuration of the shift vectors appearing in a lattice equation. In the case of quad equations, the polytope obtained as the convex hull of the shift vectors appearing in an equation, i.e., $\operatorname{conv}\{(0,0),(-1,0),(0,-1),(-1,-1)\}$, is a unit square. The polytope has two edges attached to the origin $(0,0)$, but neither of them has a shift vector in the middle. This property played an essential role in the proofs of Lemmas 4.9 and 4.10. Therefore, to prove these lemmas for general lattice equations would require some knowledge of convex geometry.
@@ Appendix A Essential choices of domain
Appendix B Substitution and derivatives
In this appendix, let $K$ be an arbitrary field of characteristic 00 and consider the rational function field $K(z,w)$. We allow $\infty$ as the result of substitution throughout this paper. This is why we need some unusual definitions and lemmas about substitution and derivatives.
@@ Appendix B Substitution and derivatives
Next, we show the converse. Let $\partial_{w}f(z^{*},w)=0$. If $f(z^{*},w)=\infty$, then it is obviously independent of $w$. Therefore, we may assume that $f(z^{*},w)\neq\infty$. In this case, usual properties of the derivative hold and thus $f(z^{*},w)$ does not depend on $w$. If $\partial_{w}\frac{1}{f(z^{*},w)}=0$, one can show in the same way that $g(z^{*},w)$ does not depend on $w$ where $g(z,w)=\frac{1}{f(z,w)}$. ∎